OpenStudy (anonymous):

find the first term of the arithmetic sequence an = 48, Sn = 546, n = 26

OpenStudy (anonymous):

an= a1+ (n-1)·d --> 48=a1+25·d Sn= (a1+an)·n/2 = 546 --> (a1+48)·13= 546 --> a1+48=42

OpenStudy (anonymous):

so its -6??

mathslover (mathslover):

yes ..

OpenStudy (anonymous):

thanks

mathslover (mathslover):

pleasure